Electromagnetic calorimeters based on scintillating lead tungstate crystals for experiments at Jefferson Lab^{}

A.Asaturyan^{a}, F.Barbosa^{c}, V.Berdnikov^{b}, J.Crafts^{b,c}, H.Egiyan^{c}, L.Gan^{f}, A.Gasparian^{g}, K.Harding^{c}, T.Horn^{b}, V.Kakoyan^{a}, H.Mkrtchyan^{a}, Z.Papandreou^{e}, V. Popov^{c}, S.Taylor^{c}, N.Sandoval^{c}, A.Somov^{c,∗}, S.Somov^{d}, A. Smith^{h}, C. Stanislav^{c}, H.

Voskanyan^{a}, T. Whitlatch^{c}, S. Worthington^{c}

aA. I. Alikhanian National Science Laboratory (Yerevan Physics Institute), 0036 Yerevan, Armenia

bCatholic University of America, Washington, DC 20064, USA cThomas Jefferson National Accelerator Facility, Newport News, VA 23606, USA

dNational Research Nuclear University MEPhI, Moscow 115409, Russia

eUniversity of Regina, Regina, Saskatchewan, Canada S4S 0A2 fUniversity of North Carolina at Wilmington, Wilmington, NC 28403, USA

gNorth Carolina A&T State University, Greensboro, NC 27411, USA hDuke University, Durham, NC 27708, USA

Abstract

A new electromagnetic calorimeter consisting of 140 lead tungstate (PbWO_{4}) scintillating crystals was constructed for the PrimEx η experiment at Jefferson lab. The calorimeter was integrated into the data acquisition and trigger systems of the GlueX detector and used in the experiment to reconstruct Compton scattering events. The experiment started collecting data in the spring of 2019 and acquired about 30% of the required statistics. The calorimeter is a prototype for two PbWO_{4}-based detectors: the Neutral Particle Spectrometer (NPS) and the lead tungstate insert of the forward calorimeter (FCAL) of the GlueX detector. The article presents the design and performance of the Compton calorimeter and gives a brief overview of the FCAL and NPS projects.

Keywords: Electromagnetic calorimeter, Lead tungstate scintillator

1 1. Introduction

2 Electromagnetic calorimeters based on PbWO_{4} scintillating 3 crystals have a widespread application in experiments at differ-4 ent accelerator facilities such as CERN, FNAL, GSI, and Jef-5 ferson Lab (JLab). The small radiation length (L_{R} = 0.89 cm) 6 and Molière radius (R_{M} = 2.19 cm) of PbWO_{4} allows to 7 build high-granularity detectors with a good spatial separation 8 and energy resolution of reconstructed electromagnetic show-9 ers, which makes these crystals the material of choice in many 10 of these applications. 11 Two electromagnetic calorimeters are currently under con-12 struction in experimental Hall D and Hall C at Jefferson Lab, 13 both using rectangular 2.05 cm × 2.05 cm × 20 cm PbWO_{4} 14 scintillating modules. The inner part of the forward lead glass 15 calorimeter of the GlueX detector [1] in Hall D will be upgraded 16 with these high-granularity, high-resolution crystals. This up-17 grade is required by the physics program with the GlueX de-18 tector, specifically the new experiment to study rare decays of 19 η mesons [2]. The size of the insert will tentatively consist of 20 2496 lead tungstate modules. The Neutral Particle Spectrome-21 ter [3] in experimental Hall C consists of a PbWO_{4} electromag-22 netic calorimeter preceded by a sweeping magnet. The NPS is 23 required by Hall C’s precision cross section measurement pro-24 gram with neutral final states [4–9]. Such precision measure-25 ments of small cross sections play a central role in studies of 26 transverse spatial and momentum hadron structure. The NPS 27 detector consists of 1080 PbWO_{4} crystals arranged in a 30 × 36 28 array. Lead tungstate crystals for both detectors were procured 29 from two vendors: Shanghai Institute of Ceramics (SICCAS) 30 in China and CRYTUR in the Czech Republic. The quality of 31 recently produced PbWO_{4} scintillators has been studied in de-32 tail by the NPS and EIC eRD1 collaborations and is described 33 in Ref. [10]. PbWO_{4} crystals are also being considered for 34 an electromagnetic calorimeter of the future Electron-Ion Col-35 lider [11]. 36 In this article we describe the design and construction of 37 a calorimeter prototype composed of 140 SICCAS crystals, 38 which served as the Compton Calorimeter (CCAL) in the 39 PrimEx η experiment [12] with the GlueX detector in the spring 40 of 2019. The CCAL was subsequently used during a few short 41 GlueX physics runs at high luminosity in order to study rates 42 and operating conditions expected for the FCAL lead-tungstate 43 insert. Experience gained during fabrication and operation of 44 the CCAL was critical for finalizing the design of the FCAL 45 insert and also helped further optimize the NPS calorimeter. 46 This article is organized as follows: we will present the

Notice: Authored by Jefferson Science Associates, LLC under U.S. DOE Contract No. DE-AC05-06OR23177. The U.S. Government retains a nonexclusive, paid-up, irrevocable, world-wide license to publish or reproduce this manuscript for U.S. Government purposes.

∗Corresponding author. Tel.: +1 757 269 5553; fax: +1 757 269 6331.

Email address: somov@jlab.org (A.Somov)

Preprint submitted to Elsevier March 30, 2021

Figure 1: Schematic layout of the GlueX detector (not to scale). Numbers represent the following detector components: solenoid magnet (1), barrel calorimeter (2), central drift chamber (3), forward drift chambers (4), time-of-flight wall (5).

47 PrimEx η experiment and performance of the CCAL in Sec-48 tion 2 and Section 3, and will briefly describe the FCAL and 49 NPS projects in Sections 4 and 5.

50 2. PrimEx η experiment with the GlueX detector

51 The GlueX detector [1] was designed to perform experi-52 ments using a photon beam. Beam photons are produced via 53 the bremsstrahlung process by electrons, provided by the JLab 54 electron accelerator facility, incident on a thin radiator. The en-55 ergy of a beam photon is determined by detecting a scattered 56 electron after radiating the photon with a typical precision of 57 0.2%. The electron is deflected in a 6 m long dipole mag-58 net operated at a field of 1.8 T and registered in the so-called 59 tagging scintillator counters. Each counter corresponds to the 60 specific energy of the reconstructed lepton. The photon beam 61 propogates toward the GlueX target. A schematic view of the 62 GlueX detector is illustrated in Fig. 1^{1}. 63 The physics goal of the PrimEx η experiment is to perform a 64 precision measurement of the η → γγ decay width. The mea-65 surement will provide an important test of quantum chromody-66 namics symmetries and is essential for the determination of fun-67 damental properties such as the ratios of the light quark masses 68 and the η-η′ mixing angle. The decay width will be extracted 69 from the measurement of the photoproduction cross section of

Figure 2: Schematic layout of the Compton calorimeter.

70 η mesons in the Coulomb field of a nucleus, which is known 71 as the Primakoff effect. The η mesons will be reconstructed by 72 detecting two decay photons in the forward calorimeter of the 73 GlueX detector. 74 The cross section will be normalized using the Compton scat-75 tering process, which will also be used to monitor the luminos-76 ity and control the detector stability during data taking. Elec-77 trons and photons originating from Compton events in the target 78 are produced at small angles, typically outside the acceptance 79 of the FCAL. In order to improve the reconstruction of particles 80 in the forward direction, we built a small Compton calorimeter 81 consisting of 140 lead tungstate scintillating crystals and posi-82 tioned it about 6 m downstream from the FCAL as shown in in 83 Fig. 1. The CCAL covers the angular range between 0.19^{∘} and 84 0.47^{∘}. 85 The PrimEx η experiment started collecting data in the spring 86 of 2019 and has acquired 30% of the required statistics. Dur-87 ing the experiment, the magnetic field of the solenoid magnet 88 was switched off in order to allow reconstruction of Compton 89 events. The photon flux was about 5 · 10^{6} γ/sec (about five 90 times lower than the nominal GlueX flux) in the beam energy 91 range of interest between 9.5 GeV and 11.6 GeV.

92 3. Compton calorimeter of the PrimEx η experiment

93 3.1. Calorimeter design

94 The calorimeter design is shown in Fig. 2. The CCAL com-95 prises an array of 12 × 12 lead tungstate modules with a 2 × 2 96 hole in the middle for the passage of the photon beam. The 97 modules are positioned inside a light tight box. A tungsten ab-98 sorber is placed in front of the innermost layer closest to the 99 beamline to provide protection from the high rate of particles 100 predominantly originating from electromagnetic interactions. 101 The light yield from PbWO_{4} crystals depends on tempera-102 ture with a typical temperature coefficient of 2%/^{∘}C at room 103 temperature. Maintaining constant temperature is essential for 104 the calorimeter operation. The calorimeter modules are sur-105 rounded by four copper plates with built-in pipes to circulate a 106 cooling liquid and provide temperature stabilization. Foam in-107 sulation surrounds the detector box. The temperature was mon-108 itored and recorded during the experiment by five thermocou-109 ples attached to different points of the PbWO_{4} module assem-110 bly. During the experiment the temperature was maintained at 111 17^{∘} ± 0.2^{∘}C. The typical heat released by the photomultiplier 112 tube (PMT) dividers was equivalent to about 30 Watts. In or-113 der to prevent condensation, a nitrogen purge was applied. Two 114 fans with a water-based cooling system were installed on the 115 top of the crystal assembly to improve nitrogen circulation and 116 heat dissipation from the PMT dividers. The detector was po-117 sitioned on a platform, which allowed to move it in the vertical 118 and horizontal directions, perpendicular to the beam. The plat-119 form was remotely controlled and provided a position accuracy 120 of about 200 μm. During detector calibration each module was 121 moved into the beam.

1Not shown on this plot is the DIRC detector, which was installed after the PrimEx η experiment and is used for the particle identification in the forward direction.

122 3.2. Module design

123 The design of the PbWO_{4} module is based on the HyCal 124 calorimeter, which was used in several experiments in Jeffer-125 son Lab Hall B [13]. An assembled calorimeter module is pre-126 sented in Fig. 3. Each lead tungstate crystal is wrapped with a 127 60 μm polymer Enhanced Specular Reflector film (ESR) man-128 ufactured by 3M^{TM}, which allows 98.5% reflectivity across the 129 visible spectrum. In order to improve optical isolation of each 130 module from its neighbors, each crystal is wrapped with a layer 131 of 25 μm thick Tedlar. The PMT is located inside a G-10 fiber-132 glass housing at the rear end of the crystal. Two flanges are 133 positioned at the crystal and housing ends and are connected 134 together using 25 μm brass straps, which are brazed to the sides 135 of the flanges. Four set screws are pressed to the PMT housing 136 flange to generate tension in the straps and hold the assembly 137 together. Light from the crystal is detected using a ten-stage 138 Hamamatsu PMT 4125, which is inserted into the housing and 139 is coupled to the crystal using optical grease (EJ-550). The 140 PMT diameter is 19 mm. The PMT is pushed towards the crys-141 tal by using a G-10 retaining plate attached to the back of the 142 PMT and four tension screws applied to the PMT flange. The 143 PMT is instrumented with a high-voltage divider and ampli-144 fier positioned on the same printed circuit board attached to the 145 PMT socket.

146 3.3. Electronics

147 The PMT of each calorimeter module was equipped with an 148 active base prototype [14], which was designed for the Neutral 149 Particle Spectrometer in experimental Hall C. The base com-150 bines a voltage divider and an amplifier powered by the current 151 flowing through the divider. The active base allows the oper-152 ation of the PMT at lower voltage and consequently at lower

Figure 3: Calorimeter module showing main components: the PbWO_{4} crystal, PMT housing, PMT divider, and signal and HV cables.

153 anode current, which improves the detector rate capability and 154 prolongs the PMT’s life. The original Hamamatsu divider for 155 this type of PMT was modified by adding two bipolar transis-156 tors on the last two dynodes, which provides gain stabilization 157 at high rate. The active base has a relatively large amplification 158 of about a factor of 24 due to the large PMT count rate pre-159 dicted by Monte Carlo simulation of the NPS detector. Large 160 amplification was not needed for the planned run conditions 161 of the PrimEx η experiment. However, we subsequently used 162 CCAL in GlueX runs at significantly larger luminosity in or-163 der to study run conditions of the FCAL lead tungstate insert, 164 where the amplifier will be required. This will be discussed 165 in Section 4.0.3. During the PrimEx run, CCAL PMTs were 166 operated at about 680 V, which produced a divider current of 167 260 μA. The high voltage for each PMT was supplied by a 168 24-channel CAEN A7236SN module positioned in a SY4527 169 mainframe.

170 Amplified PMT signals were digitized using a twelve-bit 16-171 channel flash ADCs electronics module operated at a sampling 172 rate of 250 MHz. The ADC was designed at Jefferson Lab [15] 173 and is used for the readout of several sub-detectors of the GlueX 174 detector. The Field-Programmable Gate Array (FPGA) chip 175 inside the ADC module allows the implementation of various 176 programmable data processing algorithms for the trigger and 177 readout. An example of a flash ADC signal pulse obtained 178 from a calorimeter module is shown in Fig. 4. In this example, 179 the ADC is operated in the raw readout mode, where digitized 180 amplitudes are read out for 100 samples, corresponding to the 181 400 ns read out window. During the PrimEx η experiment, the 182 ADC performed on-board integration of signal pulses, which

Figure 4: A typical flash ADC signal pulse obtained from a PbWO_{4} module.

183 amplitudes were above a threshold of 24 MeV. Amplitudes were 184 summed in a time window of 64 ns and read out from the ADC 185 module along with other parameters such as the pulse peak am-186 plitude, pulse time, and data processing quality factors. This 187 readout mode allowed to significantly reduce the data size and 188 ADC readout time, and therefore did not induce any dead time 189 in the data acquisition. 190 CCAL flash ADCs are positioned in a VXS (ANSI/VITA 191 41.0 standard) crate. VXS crates are used to host all readout 192 electronics of the GlueX experiment. In addition to the VME-193 bus used to read out data from electronics modules, the VXS is 194 instrumented with a high-speed serial bus in order to increase 195 the bandwidth to several Gb/sec and provide an interconnected 196 network between modules. The bus is used to transmit ampli-197 tudes digitized by the ADC to trigger electronics modules to 198 include the CCAL in the Level 1 trigger system of the GlueX 199 detector.

200 3.4. Light Monitoring System

201 To monitor performance of each calorimeter channel, we 202 designed an LED-based light monitoring system (LMS). The 203 LMS optics includes a blue LED, a spherical lens to correct 204 the conical dispersion of the LED, and a diffusion grating to 205 homogeneously mix the light. Light produced by the LED is 206 incident on a bundle of plastic optical fibers (Edmund Optics) 207 with a core diameter of 250 μm. Each fiber distributes light to 208 an individual calorimeter module. On the crystal end, the fiber 209 is attached to the module using a small acrylic cap glued to the 210 crystal with a hole drilled through each cap to hold the fiber 211 inside. 212 To monitor stability of the LED, we used two reference

Figure 5: Flash ADC signal amplitudes induced by the LED and the α-source in the reference PMT.

213 Hamamatsu 4125 PMTs, the same type as in the CCAL detec-214 tor. Each PMT receives light from two sources: a single fiber 215 from the LED and a YAP:Ce pulser unit, both glued to the PMT 216 face. The pulser unit consists of a 0.15 mm thick YAP:Ce scin-217 tillation crystal with a diameter of 3 mm spot activated by an 218 ^{241}Am α source. The α source is used to monitor stability of 219 the LED. The PMT was read out using a flash ADC. The high 220 voltage on each reference PMT was adjusted to have the sig-221 nals from both the LED and α source fit within the range of a 222 12-bit flash ADC corresponding to 4096 counts, as shown in 223 Fig. 5. Each LED was driven by a CAEN 1495 module, which 224 allowed to generate LED pulses with a programmable rate. The 225 LMS was integrated into the GlueX trigger system and provided 226 a special trigger type during data taking. The LMS was ex-227 tensively used during the detector commissioning and injected 228 light to the CCAL detector with a typical frequency of 100 Hz 229 continuously during the PrimEx η experiment. This LED rate 230 is similar to the trigger rate of events produced by the reference 231 α source.

232 Most LMS components were positioned inside the 233 temperature-stabilized detector box. The stability of the 234 LED system measured using the reference PMTs during the 235 entire PrimEx run was on the level of 1%. The ratio of signal 236 ADC amplitudes from the LED pulser to the α source obtained 237 during different run periods of the 48-day long PrimEx η 238 experiment is presented in Fig. 6. The ratio is normalized 239 to the data in the beginning of the experiment. Stability of 240 most CCAL modules observed using the LMS during the 241 experiment was better than 6%. We did not apply any PMT 242 gain adjustments during the experiment.

Figure 6: Ratio of signal ADC amplitudes from the LED pulser to the α-source measured by the reference PMT during different run periods of the 48-day long PrimEx η experiment. The ratio is normalized to data in the beginning of the run.

243 3.5. Calibration

244 The initial energy calibration of the CCAL was performed by 245 moving the calorimeter frame and positioning each module into 246 the photon beam during special low-intensity calibration runs. 247 The maximum rate in the module exposed to the beam did not 248 exceed 200 kHz at a threshold of 15 MeV. The energy of each 249 beam photon was determined by detecting a bremsstrahlung 250 electron using the GlueX tagging detectors described in Sec-251 tion 2. The tagging detectors cover the photon energy range 252 between 2.9 GeV and 11.4 GeV and provide the relative en-253 ergy resolution of about 0.2%. The spot size of the collimated 254 photon beam had a diameter of about 6 mm. 255 In the beginning of the calibration run, we adjusted the PMT 256 high voltage for each module in order to equalize signal pulse 257 amplitudes induced by 10 GeV beam photons. The amplitude 258 was set to 3200 ADC counts. An example of flash ADC signal 259 amplitude in the calorimeter module as a function of the beam 260 energy is presented in Fig. 7. The calibration of each module 261 was refined by reconstructing showers in the calorimeter and 262 constraining the reconstructed energy to the known beam en-263 ergy. 264 During the calibration runs, we estimated the non-uniformity 265 of the 140 CCAL modules by measuring the relative energy 266 resolution for each individual module exposed to the beam. The 267 energy resolution obtained for 6 GeV photons is presented in 268 Fig. 8. The distribution is fit to a Gaussian function. The non-269 uniformity of the modules, i.e., the spread of the distribution is 270 found to be smaller than 5%. 271 During calibration, we observed some a non-linearity of the

Figure 7: ADC signal pulse amplitude in the CCAL module as a function of the beam energy.

272 PMT active base with the large amplification factor of 24, on 273 the level of a few percent, which impacted both the pulse peak 274 and pulse integral. The base performance became linear when 275 the amplifier gain was reduced. In order to study the impact of 276 the non-linearity on the detector energy resolution, we replaced 277 the original PMT active bases for 9 CCAL modules (in the array 278 of 3 × 3 modules) with modified bases where the amplifier was 279 bypassed. After adjusting high voltages and recalibrating PMT 280 gains, we measured the energy resolution for different beam en-281 ergies. The beam was incident on the center of the middle mod-282 ule in the array. An example of the energy deposited by 10 GeV 283 photons is shown in Fig. 9. The energy resolution was obtained 284 from a fit of the energy distribution to a Crystal Ball function^{2} 285 implemented in the ROOT data analysis framework [16]. The 286 energy resolution as a function of the beam energy is shown in 287 Fig. 10. The distribution was fit to the following function:

288  \frac{σ_{E}}{E} = \frac{S}{\sqrt{E}} ⊕ \frac{N}{E} ⊕ C,  (1)

289 where S represents the stochastic term, N the electronic noise 290 and C the constant term, E is the beam energy in GeV, and 291 the symbol ⊕ indicates a quadratic sum. The fit yields: S = 292 2.63 ± 0.01%, N = 1.07 ± 0.09%, and C = 0.53 ± 0.01%. The 293 resolution was found to be about 10% better than that measured 294 with the original base with the amplifier gain of 24 ^{3}. The en-295 ergy resolution is consistent with that of the HyCal calorime-296 ter [13], which was instrumented with crystals produced by

2The function is named after the Crystal Ball collaboration.

3The linearity of the PMT active base is being currently improved; modified active bases will be installed before the new PrimEx η run in 2021.

Figure 8: Relative energy resolution of 140 PbWO_{4} modules installed on the CCAL measured with 6 GeV beam photons.

297 SICCAS in 2001 and was used in several experiments in Jef-298 ferson Lab’s experimental Hall B. The HyCal PbWO_{4} crystals 299 have the same transverse size of 2.05 cm×2.05 cm, but a smaller 300 length of 18 cm. The initial CCAL calibration performed with 301 the beam scan was fine-tuned during the PrimEx ηrun by using 302 showers of reconstructed Compton scattering candidates and 303 constraining the reconstructed energy in the event to the know 304 beam energy.

305 3.6. Performance during the PrimEx η run

306 In the PrimEx η experiment, we reconstruct Compton events 307 produced by beam photons with E_{beam} > 6 GeV. This energy 308 range is covered by the GlueX pair spectrometer [17], which 309 determines the photon flux needed for cross section measure-310 ments. An electron and photon produced in the Compton scat-311 tering process were detected by reconstructing two showers, 312 one in the FCAL and another one in CCAL. The event topology 313 of the reaction is such that the more energetic electron predom-314 inantly goes into the Compton calorimeter, while the photon 315 strikes the FCAL. In order to accept Compton events during 316 data taking and to reduce background originating from low-317 energy electromagnetic and hadronic interactions, the CCAL 318 was integrated to the Level 1 trigger system of the GlueX de-319 tector. The physics trigger was based on the total energy de-320 posited in the forward and Compton calorimeters. The GlueX 321 trigger is implemented on special-purpose programmable elec-322 tronics modules with FPGA chips. The trigger architecture is 323 described in Ref. [18]. The trigger rate as a function of the en-324 ergy threshold is presented in Fig. 11. We collected data using 325 a relatively small energy threshold of 3 GeV at a trigger rate

Figure 9: Energy distribution deposited by 10 GeV beam photons. The spectrum is fit to a Crystal Ball function.

326 of about 18 kHz. This rate did not produce any dead time in 327 the data acquisition and trigger systems. The trigger rate was 328 reproduced by a detailed Geant detector simulation. 329 The rate in the CCAL modules during the experiment is pre-330 sented in Fig. 12. In this plot, the photon beam goes through the 331 center of the hole of 2× 2 modules in the middle of the detector. 332 The rate is the largest in innermost detector layers closest to the 333 beam line. The maximum rate in the detector module was about 334 200 kHz for an energy threshold of 30 MeV, which is equivalent 335 to a signal pulse amplitude of 5 mV. Before the experiment, we 336 performed a high-rate performance study of the PMT and elec-337 tronics using a laser and an LED pulser and did not find any 338 degradation of the PMT gain in run conditions similar to the 339 PrimEx η experiment up to 3-4 MHz [19]. 340 Timing resolution of reconstructed showers is an important 341 characteristic of the detector performance. In the experiment 342 we used timing information provided by the calorimeters to 343 identify the accelerator beam bunch for which the interaction 344 occurred in the detector and therefore relate showers in the 345 calorimeters with hits in the tagging detector, from the same 346 event. A hit in the tagging detector defines the energy of the 347 beam photon. The time in the calorimeter module is provided 348 by an algorithm implemented on the programmable FPGA chip 349 of the flash ADC. The algorithm performs a search of the peak 350 of the signal pulse and determines the time from the shape of 351 the leading edge of the pulse. The times of all hits constituting 352 the CCAL shower are combined to form the shower time by 353 using an energy-weighted sum. The time difference between 354 beam photon candidates and CCAL showers originating from 355 Compton events is presented in Fig. 13. The main peak on this

Figure 10: Energy resolution as a function of the photon energy.

356 plot corresponds to beam photons and CCAL clusters produced 357 in the same accelerator bunch. Satellite peaks, separated by the 358 beam bunch period of about 4 ns, represent accidental beam 359 photons from accelerator bunches not associated with the inter-360 action in the detector. The time resolution of CCAL showers is 361 improved with the increase of the shower energy and was mea-362 sured to be about 330 ps and 140 ps for 1 GeV and 9 GeV show-363 ers, respectively. In the PrimEx η experiment, CCAL allowed 364 a clear separation of beam photons originating from different 365 beam bunches. 366 Reconstruction of electromagnetic showers in the FCAL is 367 performed using an algorithm described in Ref. [20], which is 368 a part of the standard GlueX reconstruction software. For the 369 CCAL, we implemented an algorithm originally developed for 370 the GAMS spectrometer [21], which was subsequently adopted 371 for the HyCal [13] in JLab’s experimental Hall B. The algo-372 rithm provides a good separation of overlapping showers in the 373 calorimeter by using profiles of electromagnetic showers. The 374 elasticity distribution, defined as the reconstructed energy in 375 the event minus the beam energy, is presented in Fig. 14 for 376 Compton candidates produced by beam photons in the energy 377 range between 6 GeV and 7 GeV. The solid line shows the fit 378 of this distribution to the sum of a Gaussian and a second order 379 polynomial function. The energy resolution of reconstructed 380 Compton candidates in this energy range is about 130 MeV. In 381 this plot, we subtracted background originating from accidental 382 beam photons. This background was measured using off-time 383 interactions and amounted to about 15%. The relatively small 384 background, on the level of 10%, produced by interactions of 385 the photon beam with the beamline material downstream the 386 GlueX target was measured using empty-target runs and was

Figure 11: Trigger rate as a function of the total energy deposited in the FCAL and CCAL. The arrow indicates the energy threshold used in PrimEx η production runs.

387 also excluded from the elasticity distribution in Fig. 14. The 388 CCAL allowed to clearly reconstruct Compton events in the 389 PrimEx η experiment.

390 4. Upgrade of the GlueX forward calorimeter

391 The forward calorimeter of the GlueX detector con-392 sists of 2800 lead glass modules, each with a size of 393 4 cm × 4 cm × 45 cm, and is positioned about 6 m down-394 stream of the target, as shown in Fig. 1. The FCAL covers 395 a polar angle of photons produced from the target between 1^{∘} 396 and 11^{∘} and detects showers with energies in the range of 0.1 397 - 8 GeV. The Cherenkov light produced in the module is de-398 tected by FEU-84-3 photomultiplier tubes, instrumented with 399 Cockcroft-Walton bases [22]. The typical energy resolution of 400 the FCAL is σ_{E}/E = 6.2%/ \sqrt{E} ⊕ 4.7%. 401 The future physics program with the GlueX detector in 402 Hall D will require an upgrade of the inner part of the for-403 ward calorimeter with high-granularity, high-resolution PbWO_{4} 404 crystals. The lead tungstate insert will improve the separa-405 tion of clusters in the forward direction and the energy reso-406 lution of reconstructed photons by about a factor of two. Lead 407 tungstate crystals possess better radiation hardness compared to 408 lead glass, which is important for the long term operation of the 409 detector at high luminosity. We propose to build a 1 m × 1 m 410 insert, which will require about 2496 modules. Similar to the 411 CCAL, the insert will have a beam hole of 2 × 2 modules and 412 a tungsten absorber used to cover the detector layer closest to 413 the beamline. A schematic view of the FCAL frame with the 414 installed lead tungstate insert is presented in Fig. 15. Due to the

Figure 12: Rates in the CCAL modules during PrimEx η production run. The energy threshold corresponds to 30 MeV. The beam goes through the center of the hole in the middle of the plot.

415 different size of the lead glass bars and lead tungstate crystals, 416 the lead glass modules stacked around the PbWO_{4} insert will 417 form four regions with a relative offset between modules; those 418 regions are shown in green color in this plot. 419 The PbWO_{4} module design of the FCAL insert will essen-420 tially be the same as for the CCAL, except for some small 421 modifications needed to handle the magnetic field present in 422 the FCAL region. The PMT housing made of the G-10 fiber-423 glass material will be replaced by iron housing in order to re-424 duce the magnetic field. The housing length will be increased to 425 extended the magnetic shield beyond the PMT photo cathode. 426 An acrylic optical light guide will be inserted inside the PMT 427 housing to couple the crystal and PMT. 428 The upgraded FCAL will be operated in GlueX experiments 429 using a 30 cm long liquid hydrogen target at the designed pho-430 ton flux of 5 · 10^{7} γ/sec in the energy range between 8 GeV 431 and 9 GeV. The designed luminosity is significantly larger than 432 that used in the PrimEx η experiment and was achieved after the 433 PrimEx run in the fall of 2019. In order to finalize the design of 434 the PMT electronics, it is important to understand detector rates 435 in the FCAL insert, especially in layers close to the beamline. 436 We used CCAL during high-intensity GlueX runs to study run 437 conditions for the FCAL insert.

438 4.0.1. PMT magnetic shield 439 The longitudinal (directed along the beamline) and transverse 440 (directed perpendicular to the axis of of the beamline) compo-441 nents of the magnetic field produced by the GlueX solenoid 442 magnet in the FCAL PbWO_{4} insert area vary between 40 - 50 443 Gauss and 0 - 8 Gauss, respectively. The longitudinal field is 444 the largest on the beamline, where the transverse component 445 is practically absent. We studied the PMT magnetic shield-

Figure 13: Time difference between beam photons and reconstructed CCAL showers for Compton candidates. Peaks are separated by the beam bunch period of 4 ns.

446 ing using a prototype consisting of an array of 3 × 3 PMT 447 iron housings made of AISI 1020 steel, which was positioned 448 in the middle of Helmholtz coils. Each housing had a size of 449 20.6 mm × 20.6 mm × 100 mm with a 19.9 mm round hole in 450 the middle for the PMT. This corresponds to the realistic size of 451 the magnetic shield that will be used in the calorimeter module 452 assembly. Inside the housing we inserted two layers of μ-metal 453 Co-Netic cylinders, with thicknesses of 350 μm and 50 μm, 454 separated from each other by a Kapton film. The thickest cylin-455 der was spot welded and annealed. 456 The Helmholtz coils had a diameter of about 1 m and can 457 generate a uniform magnetic field with variable strength below 458 100 Gauss. A Hall probe was inserted into the central mod-459 ule of the prototype to measure the magnetic field at different 460 Z-positions along the length of the cylinder. The field was mea-461 sured for two different orientations of the prototype with respect 462 to the magnetic field: field oriented along the PMT (longitudi-463 nal, B_{z}) and perpendicular to the PMT housing (transverse, B_{x}). 464 Field measurements are presented in Fig. 16. The PMT shield 465 significantly reduces both the longitudinal and transverse fields 466 to the level of B_{z} ∼ 1 Gauss and B_{x} ≪ 1 Gauss. The trans-467 verse field, which is well shielded, is more critical for the PMT 468 operation, as it is directed perpendicular to the electron trajec-469 tory inside the photo tube and deflects electrons, resulting in 470 the degradation of the photon detector efficiency and gain. The 471 field reaches a plateau at Z = 3 cm from the face of the hous-472 ing. We will use 3.5 cm long acrylic light guides, in order to 473 place the most sensitive to the magnetic field area of the PMT 474 between the photocathode and the last dynode (4.6 cm long) in 475 the region with the smallest magnetic field, as shown in Fig. 16.

Figure 14: Elasticity distribution of reconstructed Compton candidates.

476 We studied performance of the shielded PMT in the magnetic 477 field using an LED pulser. A blue LED with a light diffuser was 478 placed about 20 cm from the PMT housing prototype and was 479 aligned with the middle module. The PMT response was mea-480 sured for different pulse amplitudes and operational high volt-481 ages. In order to study the contributions from longitudinal and 482 transverse field components we rotated the prototype by dif-483 ferent angles. Signal amplitudes as a function of the magnetic 484 field measured in the prototype tilted by about 10 degrees are 485 presented on the left plot of Fig. 17. Amplitudes, normalized to 486 measurements without magnetic field, are shown on the bottom 487 plot. The relative degradation of the signal amplitude for the 488 maximum field in the FCAL insert region of B = 50 Gauss (B_{z} 489 = 49 Gauss and B_{x} = 8.6 Gauss) was measured to be less than

490 1%.

491 4.0.2. Light guide studies 492 Studies of the magnetic shielding demonstrated that the PMT 493 has to be positioned inside the iron housing at the distance of 494 at least 3 cm from the face of the PbW0_{4} crystal, where the 495 magnetic field is reduced and reaches the plateau. In order to 496 do this, in the FCAL insert module we decided to use a 3.5 cm 497 long acrylic cylindrical light guide with a diameter of 18.5 mm 498 between the PMT and the PbWO_{4} crystal. The light guide is 499 wrapped with reflective ESR foil and attached to the PMT with 500 Dymax 3094 UV curing glue. Optical coupling to the crystal is 501 provided by a “silicon cookie”: a 1 mm thick transparent rub-502 ber cylinder made of the room temperature vulcanized silicon 503 compound, RTV615. The silicon cookie is not glued to the light 504 guide and the crystal, so the module can be easily disassembled 505 if its PMT needs to be replaced.

Figure 15: FCAL frame with calorimeter modules installed: PbWO_{4} crystals (brown area), lead glass blocks (green). The photon beam passes through the hole in the middle of the calorimeter.

506 We compared light losses of the FCAL insert module instru-507 mented with the light guide with the CCAL module, where the 508 PMT was coupled directly to the crystal using an optical grease. 509 Light collection was measured using electrons provided by the 510 Hall D pair spectrometer (PS) [17]. The PS is used to mea-511 sure the flux of beam photons delivered to the experimental hall 512 by detecting electromagnetic electron-positron pairs produced 513 by the photons in a thin converter inserted to the beam. Lep-514 tons from the pair are deflected in a dipole magnet and detected 515 using scintillator detectors placed in the electron and positron 516 arms of the spectrometer. The energy of a lepton is detected 517 using a high-granularity PS hodoscope, which consists of 145 518 scintillating tiles and covers the energy range between 3 GeV 519 and 6 GeV. The relative light yield of the module with and with-520 out the light guide was estimated by positioning the module be-521 hind the PS and measuring signal amplitudes induced by the PS 522 electrons.

523 We first measured the ADC response in the CCAL module, 524 which was subsequently modified by adding the light guide to 525 the same PMT and crystal and was placed to the same spot of 526 the PS test setup. Results of the measurements are presented 527 in Fig. 18. The ADC amplitude of the calorimeter module is 528 presented as a function of the PS tile for the two module con-529 figurations with and without the light guide. The light guide re-530 sults in a relatively small loss of light of about 15% compared 531 with the CCAL module. We note that wrapping the light guide 532 with the reflective material is important. Losses in unwrapped 533 light guide constitute about 35%. We repeated light collection 534 measurements using two more modules and obtained consistent 535 results.

Figure 16: Magnetic field distribution inside the PMT shield housing as a function of the distance from the housing face. Plot (a) corresponds to the longitudinal field and plot (b) corresponds to the transverse field produced by the Helmholtz coils. Markers denote different field values.

536 4.0.3. Detector rate 537 The PMT anode current is one of the critical characteristics 538 that have to be considered during the design of the PMT divider. 539 Typically the anode current should be on the level of a few mi-540 cro amperes and significantly smaller than the divider current in 541 order to provide stable performance of the PMT base and pre-542 vent the long-term degradation of the PMT. The anode current 543 was measured in CCAL modules during data production runs 544 at the GlueX nominal luminosity. The special random trigger 545 was used to read out flash ADC raw data for each CCAL mod-546 ule in a time window of 400 ns, which corresponds to 100 flash 547 ADC samples. The average voltage, Ū, in the readout window 548 was determined by summing up ADC amplitudes, converted to 549 units of Volts, and normalizing the sum to the window size. The 550 anode current can be related to the average ADC voltage as

551  I = \frac{Ū}{R} · \frac{1}{G},  (2)

552 where R is the input impedance of the amplifier (∼ 50 Ω) and G 553 is the amplifier gain of 24. The typical anode current measured 554 in CCAL modules situated at different distances from the beam 555 line is presented in Fig. 19. Modules from the first CCAL layer 556 closest to the beamline and the outer most layer were not used 557 in the analysis. The inner module was covered by a tungsten 558 absorber and the outer module was obscured by the FCAL. The 559 rate in the detector is dominated by the forward-directed elec-560 tromagnetic background. The anode current is the largest in 561 the innermost layer of the detector closest to the beam line and 562 amounts to about 1.4 μA. This current is significantly smaller

Figure 17: Signal amplitudes of shielded PMT induced by an LED as a function of the magnetic field (a). Amplitudes, normalized to measurements without magnetic field (b). The PMT response was measured for different intensities of light pulse and HV settings as shown by different polymarkers.

563 than the PMT divider current of about 300 μA. We used the 564 CCAL measurements to extrapolate the current to the FCAL 565 insert. Taking the geometrical location of the FCAL and CCAL 566 into account, the largest PMT current in the insert PbWO_{4} mod-567 ule closest to the beam line was conservatively estimated to be 568 about 20 μA for a PMT base operated at 1 kV, assuming that 569 no amplifier is used. The detector rate drops rapidly with the 570 increase of the radial distance from the beamline. We are con-571 sidering to instrument PMTs in a few inner FCAL insert layers 572 with an amplifier with a gain of 5 and to omit the amplifier on 573 other modules.

574 5. Neutral Particle Spectrometer

575 The NPS is a new facility in Hall C that will allow access 576 to precision measurements of small cross sections of reactions 577 with neutral final states. The NPS consists of an electromag-578 netic calorimeter preceded by a sweeping magnet. As operated 579 in Hall C, it replaces one of the focusing spectrometers. 580 The NPS science program currently features six fully ap-581 proved experiments. E12-13-010 [4] and E12-06-114 [5] ex-582 periments will measure the Exclusive Deeply Virtual Comp-583 ton Scattering and π^{0} cross sections to the highest Q^{2} acces-584 sible at Jefferson Lab. Both experiments will provide impor-585 tant information for understanding Generalized Parton Distri-586 butions (GPDs). The E12-13-007 [6] experiment will study 587 semi-inclusive π^{0} electroproduction process and seeks to val-588 idate the factorization framework that is needed by the entire 589 12 GeV Jefferson Lab semi-inclusive deep-inelastic scattering

Figure 18: ADC amplitudes of the calorimeter module as a function of the pair spectrometer tile for two configurations: the PMT directly coupled to the PbWO_{4} crystal (circles), and the PMT coupled to the module using an optical light guide (boxes).

590 program. Measurements of Wide-Angle and Timelike Compton 591 Scattering reactions will be performed by the E12-14-003 [7] 592 and E12-17-008 [8] experiments. These measurements will 593 allow to test universality of GPDs using high-energy photon 594 beams. The NPS will also be used in the E12-14-005 [9] exper-595 iment to study exclusive production of π^{0} at large momentum 596 transfers in the process γp → π^{0}p. 597 The NPS science program requires neutral particle detection 598 over an angular range between 6 and 57.3 degrees at distances 599 of between 3 and 11 meters^{4} from the experimental target. The 600 experiments will use a high-intensity beam of electrons with 601 the energies of 6.6, 8.8, and 11 GeV, and a typical luminos-602 ity of ∼ 10^{38} cm^{−2}s^{−1} as well as a secondary beam of photons 603 incident on a liquid hydrogen target. A vertical-bend sweeping 604 magnet with integrated field strength of 0.3 Tm will be installed 605 in front of the spectrometer in order to suppress and eliminate 606 background of charged particle tracks originating from the tar-607 get. The photon detection is the limiting factor of the exper-608 iments. Exclusivity of the reaction is ensured by the missing 609 mass technique and the missing-mass resolution is dominated 610 by the energy resolution of the calorimeter. The calorimeter is 611 anticipated to provide the spacial resolution of 2-3 mm and the 612 energy resolution of about 2%/ \sqrt{E}. The NPS consists of 1080 613 PbWO_{4} crystals that form an array of 30 × 36 modules. Simi-614 larly to the FCAL insert in Hall D, the NPS will be built from 615 the crystals of the same size, and instrumented with the same 616 type of PMTs and readout electronics. The details of the me-617 chanical assembly and commissioning of the NPS are currently

Figure 19: Typical PMT anode current of CCAL modules positioned at different distances from the beamline. Circles correspond to the nominal GlueX luminosity, boxes correspond to 60% of the nominal luminosity.

618 under development and will be described in a forthcoming pub-619 lication. 620 The radiation hardness and good optical quality of lead 621 tungstate crystals are critical for the NPS calorimeter. The NPS 622 collaboration, in a synergistic effort with the EIC eRD1 con-623 sortium, has characterized to date over 1200 PbWO_{4} crystals 624 produced by CRYTUR and SICCAS from 2014 to the present. 625 The results of these studies have been published in Ref. [10]. 626 CRYTUR crystal samples were found to have greater overall 627 uniformity in transmittance and light yield, and better radiation 628 hardness. Of the samples characterized by the NPS collabora-629 tion 140 SICCAS crystals have been used in the CCAL detector.

630 6. Summary

631 We described the design and performance of the Compton 632 calorimeter, which was constructed using 140 lead tungstate 633 PbWO_{4} crystals recently produced by SICCAS. The calorime-634 ter was successfully used in the PrimEx η experiment in spring 635 of 2019 for reconstruction of Compton scattering events. The 636 CCAL served as a prototype for two large-scale electromag-637 netic calorimeters based on the PbWO_{4} crystals: the lead 638 tungstate insert of the forward calorimeter of the GlueX detec-639 tor and the neutral particle spectrometer. Experience gained 640 during construction and operation of the CCAL provided im-641 portant information for finalizing the design of FCAL PbWO_{4} 642 modules and PMT dividers and also served to further optimize 643 the NPS calorimeter. We presented the design of the FCAL lead 644 tungstate insert and gave an overview of the NPS project.

4The minimum NPS angle at 3m is 8.5 degrees; at 4m it is 6 degrees.

645 7. Acknowledgments

646 This work was supported by the Department of Energy. Jef-647 ferson Science Associates, LLC operated Thomas Jefferson 648 National Accelerator Facility for the United States Depart-649 ment of Energy under contract DE-AC05-06OR23177. This 650 work was supported in part by NSF grants PHY1714133 and 651 PHY2012430. We thank the NPS collaboration/project for pro-652 viding PbWO_{4} crystals and PMTs used in the construction of 653 the CCAL.

654 References

655 [1] S. Adhikari, et al., Nucl. Instrum. Meth. A 987, 164807 (2021). 656 [2] JLab Experiment E12-12-002A, Eta Decays with Emphasis on Rare Neu-657 tral Modes:The JLab Eta Factory (JEF) Experiment, https://www.

658  jlab.org/exp_prog/proposals/14/PR12-14-004.pdf.

659 [3] T. Horn, A PbWO_{4}-based Neutral Particle Spectrometer in Hall C at 12 660 GeV JLab, J. Phys. Conf. Ser. 587 (2015) no.1, 012048. 661 [4] JLab experiment E12-13-010, Exclusive Deeply Virtual Compton and 662 Neutral Pion Cross-Section Measurements in Hall C, https://www.

663  jlab.org/exp_prog/proposals/13/PR12-13-010.pdf.

664 [5] JLab experiment E12-06-114, Measurements of the Electron-Helicity 665 Dependent Cross Sections of Deeply Virtual Compton Scattering with

666  CEBAF at 12 GeV, https://www.jlab.org/exp_prog/proposals/
667  06/PR12-06-114.pdf.

668 [6] JLab experiment E12-13-007, Measurement of SemiInclusive π^{0} Pro-669 duction as Validation of Factorization, https://www.jlab.org/exp_

670  prog/proposals/13/PR12-13-007.pdf.

671 [7] JLab experiment E12-14-003, Wide-angle Compton Scattering at 8

672  and 10 GeV Photon Energies, https://www.jlab.org/exp_prog/
673  proposals/14/PR12-14-003.pdf.

674 [8] JLab experiment E12-17-008, Polarization Observables in Wide-Angle 675 Compton Scattering at large s, t, and u, https://www.jlab.org/exp_

676  prog/proposals/17/PR12-17-008.pdf.

677 [9] JLab experiment E12-14-005, Wide Angle, Exclusive Photoproduction

678  of π^{0} Mesons, https://www.jlab.org/exp_prog/proposals/14/
679  PR12-14-005.pdf.

680 [10] T. Horn, et al., Nucl. Instrum. Meth. A 956, 163375 (2020). 681 [11] R. Abdul Khalek, et al., Science Requirements and Detector Concepts 682 for the Electron-Ion Collider: EIC Yellow Report, arXiv:2103.05419,

683  https://arxiv.org/pdf/2103.05419.pdf.

684 [12] JLab Experiment E12-10-011, A Precision Measurement of the η Radia-685 tive Decay Width via the Primakoff Effect, https://www.jlab.org/

686  exp_prog/proposals/10/PR12-10-011.pdf.

687 [13] M. Kubantsev et al., AIP Conf. Proc. 867, no.1, 51-58 (2006). A. Gaspar-688 ian, Proceedings of the 11th International Conference on Calorimetry in 689 High-Energy Physics, 109-115 (2004). 690 [14] V. Popov and H. Mkrtchyan et al., Proceedings of the IEEE conference, 691 California, 2012. 692 [15] F. Barbosa et al., Proceedings of IEEE Nuclear Science Symposium, 693 Hawaii, USA (2007). 694 [16] R. Brun and F. Rademakers, Nucl. Instrum. Meth. A 389 (1997), 81-86. 695 [17] F. Barbosa, et al., Nucl. Instrum. Meth. A 795, 376-380 (2015). 696 [18] A. Somov, AIP Conf. Proc. 1560, no.1, 700-702 (2013). 697 [19] F. Barbosa, et al., “Characterization of the NPS and CCAL readout,” 698 GlueX-doc-3272, Jefferson Lab, (2017), https://halldweb.jlab.

699  org/doc-public/DocDB/ShowDocument?docid=3272.

700 [20] R. Jones, et al., Nucl. Instrum. Meth. A566, 366374, (2006). 701 [21] A. Lednev, Preprint IHEP 93-153, Protvino (1993). F. Binon, et al., Nucl. 702 Instrum. Meth. A 248, (1986). 703 [22] A. Brunner, et al., et al., Nucl.Instrum. Meth. A414 (1998).